Evaluate teh logarthm.
1. log^5 1/25
2. log^3 9
3. Write the equation log^12 8=3/5

Answer by drk(1908)

Evaluate the logarithms:
--
1. log^5 1/25
(Log(1/25))^5
5*log(1/25)
5(log(1) - log(25))
5(0-log(5^2)
-10log(5)
--
2. log^3 9
(log(9))^3
3log(9)
3log(3^2)
6log(3)
--
3. Write the equation log^12 8=3/5
rewrite using power of b for base as
b^(3/5) = 8^12.